# Trial Length, Pricing, and Rationally Inattentive Customers

### v2

@@ 2.3 Marketing Strategy and Subscription Models
Our work also connects to the extensive literature on customer relationship management (CRM) and customer inertia. It is well-documented that acquiring a new customer is often more expensive than retaining an existing one, leading firms to focus on maximizing customer lifetime value (CLV) (Gupta et al., 2004). Firms use various strategies to increase switching costs and lock in their customer base. While much of this literature focuses on contractual or technological lock-in (Farrell and Shapiro, 1989), our model provides a specific, micro-founded mechanism for a form of psychological lock-in: the inertia that arises from the cognitive effort required to terminate a relationship. The firm’s choice of trial length can be seen as a strategic decision to set the level of this "cognitive switching cost." This adds a new dimension to the study of customer retention, suggesting that retention can be "engineered" not just through loyalty programs but through the careful design of contractual friction.
The context of our model, digital subscriptions, is central to the modern marketing, MIS, and platform economics literature. Research in this area has explored various facets of the freemium and subscription models, including conversion from free to paid tiers (Kumar, 2014), the impact of network effects (Sundararajan, 2004), and the role of user engagement in retention. Our work contributes to this stream by focusing on a previously under-explored aspect of the model: the automatic renewal feature and the role of trial length in leveraging consumer inattention. This connects to empirical findings on the power of defaults and status quo bias in digital settings (Johnson and Goldstein, 2003). Our model provides a rational, choice-theoretic explanation for why such defaults are so powerful: overcoming them requires costly cognitive effort.
@@
Finally, our analysis of "click-to-cancel" regulations and other policy interventions speaks to the growing literature on "dark patterns" in user interface design (Luguri and Strahilevitz, 2021). These are design choices that steer users toward outcomes that benefit the firm but not necessarily the user. A long trial period paired with a complex, multi-step cancellation process can be viewed as a contractual dark pattern. Our model provides a formal economic framework for analyzing the incentives behind such practices and for evaluating the market-wide consequences of policies designed to curb them. By formalizing the trade-offs involved, we can move beyond a purely descriptive account of these practices to a predictive analysis of how firms will respond to regulation, providing a valuable tool for both managers and policymakers in the digital economy.
@@ 3.1 Primitives and Timeline
We consider a market with a unit mass of consumers, each characterized by a per-period valuation $v$ for the platform’s service. This valuation is the consumer’s private information and is drawn from a cumulative distribution function $F(v)$ with continuous density $f(v)>0$ on the support $[0,1]$. A crucial modeling choice, which represents a clear departure from the dominant literature on trial offers, is that consumers know their valuation $v$ from the outset. This assumption allows us to cleanly isolate the role of attention from the role of learning. In a learning model (Villas-Boas, 2006), the trial period’s primary function is to allow consumers to resolve uncertainty about $v$. By assuming $v$ is known, we create a setting where, from a purely informational perspective, no trial period is necessary. The existence of a non-zero optimal trial length in our model will therefore be attributable solely to attentional frictions, providing a complementary and previously unformalized rationale for this ubiquitous marketing tool.
@@
TimeStage 0Stage 1Stage 2Stage 3 Firm commits to $(T,P)$ Consumers observe $v$, decide to sign up Trial period of length $T$ Automatic renewal at price $P$ Case $v\geq P$ Renews: payoff $v-P$, firm revenue $P$ Case $v<P$ Try to cancel: succeeds w.p. $q^{*}(P,\lambda(T))$, fails w.p. $1-q^{*}$
@@ 2. MA-2
    Information Cost (Shannon): The cognitive cost to achieve a cancellation probability of $q\in(0,1)$ is given by $C(q;\lambda)=\frac{1}{\lambda}H(q)$, where $H(q)=q\ln q+(1-q)\ln(1-q)$ is the Shannon entropy function and $\lambda>0$ is the consumer’s attention sensitivity. This widely used formulation, grounded in information theory and pioneered in economics by Sims (2003), provides a rigorous micro-foundation for costly information processing. It has been successfully applied in numerous contexts, from macroeconomics to marketing, and its key feature is the convexity of the cost function (Matějka and McKay, 2015). This convexity captures the intuitive idea of increasing marginal costs to achieving perfect certainty: it is relatively easy to improve one’s chances of remembering from 50% to 60%, but it requires immense effort to improve from 99% to 100%.
@@ 3. MA-3
    Memory Decay: The consumer’s attention sensitivity $\lambda$ is a decreasing function of the trial length $T$, given by the specific functional form $\lambda(T)=\frac{\lambda_{0}}{1+\beta T}$, with $\lambda_{0}>0$ and $\beta>0$. Here, $\lambda_{0}$ represents the consumer’s baseline attention for an immediate task, which can be seen as an individual trait (e.g., being naturally organized) or as being influenced by the technological environment (e.g., access to sophisticated calendar apps). The parameter $\beta$ captures the rate of attention decay over time. This assumption is grounded in the psychology of prospective memory, which documents that the ability to remember to perform a delayed intention decays over time (Einstein1990). The hyperbolic form is chosen for its tractability and its ability to elegantly nest the standard full-attention model (when $\beta=0$). When $\beta>0$, the trial length $T$ becomes a strategic instrument for the firm to directly and predictably influence the cognitive friction faced by its customers.
@@ 3.3 Refinement Assumptions for Tractability
Assumption 1 (A1: Iso-elastic Happy-Subscriber Demand).
@@
Streaming services provide a compelling illustration of isoelastic tails. For example, after Netflix’s January 2022 price hike from \$13.99 to \$15.49 (an 11% increase), its subscriber churn rate increased from 2.3% to 3.3%, implying a demand elasticity of approximately 0.4.111Data from Antenna Research, as reported in industry analyses of Netflix’s Q1-2022 performance. Similarly, Spotify’s Premium tiers have exhibited monthly churn in the 1.5% to 2.0% range across multiple price points and markets, consistent with an inelastic demand response. [^2] See, for example, Inderes Equity Research, “Spotify: churn and pricing” (2023). With $\varepsilon<1$ in these prominent cases, our model’s prediction that policies like "click-to-cancel" will raise renewal prices modestly while sharply shortening trial lengths becomes directly relevant.
@@
The valuation distribution $F$ has a positive mass of low-value users ($F(\underline{v})>0$ for some $\underline{v}>0$) and a weakly increasing hazard rate, $h(v)=f(v)/(1-F(v))$, on the interval $[0,1]$.
@@
This assumption serves two crucial roles. First, the existence of a low-value segment ensures that there is always a pool of consumers who need to cancel, making the cancellation problem economically relevant. It also ensures that the consumer participation constraint will eventually bind as the trial length grows and the expected costs of managing the subscription become prohibitive for these low-value types. Second, the Increasing Failure Rate (IFR) property is a standard regularity condition in contract theory and industrial organization that guarantees the firm’s profit function is well-behaved (specifically, concave in price), guaranteeing a unique optimal solution (Lariviere and Porteus, 2006). It ensures that the marginal revenue from "happy" subscribers diminishes sufficiently quickly, which is necessary for a well-defined optimum. This assumption is weaker than assuming a specific parametric family; it requires only that the “conditional dropout” rate $h(v)$ does not decline.
@@
3.4 Information-Cost Paradigm
To model the cancellation decision of a consumer with $v<P$, we employ the rational inattention framework. The consumer faces a binary choice: ‘Cancel‘ or ‘Do Not Cancel‘. The ‘Cancel‘ action is optimal and yields a payoff of 0. The ‘Do Not Cancel‘ action is a mistake that results in a payment of $P$. The consumer can reduce the probability of making this mistake by devoting cognitive resources (e.g., setting reminders, mentally rehearsing the task) to monitoring the deadline. Let $q$ be the probability that the consumer successfully remembers to cancel. Achieving a high value of $q$ is cognitively costly. Following the canonical approach of Matějka and McKay (2015), we model this cost using a function based on Shannon entropy.
Definition 1 (Entropy Function).
The entropy-related function that forms the basis of the information cost is defined as:
$H(q)=q\ln q+(1-q)\ln(1-q),\qquad q\in(0,1).$
By convention, $00\ln 0=$ 0. This function measures the uncertainty inherent in a binary outcome with probability $q$. It is strictly convex on its domain, reaching a minimum (maximum uncertainty) at $q=0.5$. Its first and second derivatives are:
$H^{\prime}(q)=\ln\!\left(\frac{q}{1-q}\right),\qquad H^{\prime\prime}(q)=\frac{1}{q(1-q)}>0.$
The convexity of $H(q)$ captures the idea of increasing marginal cost of attention: it is relatively easy to move from a 50% chance of success to a 60% chance, but it requires immense effort to improve from a 98% chance to a 99% chance. This reflects the real-world difficulty of eliminating all possibility of error.
Definition 2 (Shannon Information Cost).
The cognitive cost for a consumer with attention sensitivity $\lambda$ to implement a monitoring strategy that yields a cancellation success probability of $q$ is:
$C(q;\lambda)=\frac{1}{\lambda}\,H(q).$
This formulation posits that the utility cost of attention is proportional to the amount of information processed (related to the reduction in uncertainty), where $1/\lambda$ is the unit price of information processing. A higher $\lambda$ corresponds to a lower marginal utility cost of processing information, making it "cheaper" for the consumer to be attentive.
@@ 3.5 Consumer Monitoring Problem
A consumer with valuation $v<P$ faces a trade-off. They can accept the risk of paying $P$ for a service they do not want, or they can exert costly mental effort to increase their chance of remembering to cancel. This decision is at the heart of the rational inattention framework: attention is a scarce resource to be allocated optimally. The consumer chooses a reminder probability $q\in[0,1]$ to solve the following cost-minimization problem:
$\min_{q\in[0,1]}\quad\underbrace{\bigl{(}1-q\bigr{)}P}_{\text{Expected Monetary Loss}}\;+\;\underbrace{\frac{1}{\lambda(T)}\,H(q)}_{\text{Cognitive Cost}}.$
This objective function transparently captures the core tension. The first term is the expected financial loss from inattention, which decreases as the probability of remembering, $q$, increases. The second term is the cognitive cost of achieving that probability, which increases in $q$. The consumer’s optimal choice, $q^{*}$, will balance these two competing costs at the margin.
Lemma 1 (Optimal Monitoring).
Equation (3.5) has a unique minimizer given by the logistic function:
$q^{*}(P,\lambda)=\frac{1}{1+\exp(-\lambda P)}.$
Moreover, the cancellation probability $q^{*}$ is increasing in price $P$ and attention $\lambda$, but decreasing in trial length $T$:
$\frac{\partial q^{*}}{\partial P}=\lambda q^{*}(1-q^{*})>0,\quad\frac{\partial q^{*}}{\partial\lambda}=P\,q^{*}(1-q^{*})>0,\quad\frac{\partial q^{*}}{\partial T}=-\beta\lambda_{0}P\frac{q^{*}(1-q^{*})}{(1+\beta T)^{2}}<0.$
Proof.
The derivation follows from the first-order condition of the convex minimization problem. The comparative statics are obtained by direct differentiation, with the derivative with respect to $T$ following from the chain rule. ∎
The results of this lemma are highly intuitive. Consumers will try harder to remember to cancel (i.e., choose a higher $q^{*}$) when the financial stakes are higher (a larger $P$) and when paying attention is cognitively cheaper (a higher $\lambda$). The result that $q^{*}$ decreases with trial length $T$ is a direct consequence of our memory decay assumption (MA-3) and forms the central behavioral mechanism of our model. The logistic functional form is a hallmark of the Shannon entropy cost structure and provides an analytically tractable foundation for the rest of our analysis.

### v3

@@ 2.3 Digital Economy, Subscription Models, and Customer Management
Our work also connects to the extensive literature on customer relationship management (CRM) and customer inertia. It is well-documented that acquiring a new customer is often more expensive than retaining an existing one, leading firms to focus on maximizing customer lifetime value (CLV) (Gupta et al., 2004; Venkatesan and Kumar, 2004; Fader and Toms, 2018). Existing literature has extensively studied switching costs and customer lock-in strategies (Burnham et al., 2003; Jones et al., 2007; Dey et al., 2013). While much of this literature focuses on contractual or technological lock-in (Farrell and Shapiro, 1989), our model provides a specific, micro-founded mechanism for psychological lock-in: the inertia arising from the cognitive effort required to terminate a relationship. The strategic implications extend to subscription pricing models more broadly. Danaher (2002) study how different pricing structures for usage and access influence users’ usage and retention, while Tian and Feinberg (2020) show how firms can use menu design to encourage particular subscription choices. Our model contributes to this literature by focusing on temporal aspects of subscription design, specifically how trial length can be strategically chosen to leverage consumer inattention.
@@
Finally, our analysis of recently proposed "click-to-cancel" regulations and other policy interventions speaks to the growing literature on "dark patterns" in user interface design (Luguri and Strahilevitz, 2021; Mathur et al., 2019). These are design choices that nudge users toward outcomes that benefit the firm, but not necessarily the user. A long trial period paired with a complex, multi-step cancellation process can be viewed as a contractual dark pattern. Our model provides a formal economic framework for analyzing the incentives behind such practices and for evaluating the market-wide consequences of policies designed to counter them, enabling policymakers to make informed trade-off decisions.
By formalizing the trade-offs involved, we move beyond purely descriptive accounts of these practices to predictive analysis of how firms will respond to regulation. This provides a valuable tool for both managers seeking to understand competitive dynamics in subscription markets and policymakers evaluating the welfare implications of various regulatory interventions in the digital economy.
@@ 3.1 Primitives and Timeline
We consider a market with a unit mass of consumers, each characterized by a per-period valuation $v$ for the platform’s service. This valuation is the consumer’s private information and is drawn from a cumulative distribution function $F(v)$ with continuous density $f(v)>0$ on the support $[0,1]$. An important modeling choice we make, a departure from the prior literature on trial offers, is that consumers know their product valuation $v$ from the outset. This assumption allows us to cleanly isolate the role of attention from the role of learning. In a learning model (Iyengar et al., 2007), the trial period’s primary function is to allow consumers to resolve uncertainty about $v$. By assuming $v$ is known, we create a setting where, from a purely informational perspective, no trial period is necessary. The existence of a non-zero optimal trial length in our model will therefore be attributable solely to attentional frictions, providing a complementary and so far overlooked rationale for this ubiquitous product marketing tool.
Table 1: Summary of Model Notation
Symbol Type Definition
Consumer Primitives
$v$ Parameter Consumer’s per-period valuation for the service, $v\sim F$ on $[0,1]$
$F(v),f(v)$ Parameter CDF and PDF of consumer valuations
$\tau_{0}$ Parameter Baseline attention sensitivity (at $T=0$)
$\beta$ Parameter Rate of attention decay over time
Firm / Contract Primitives
$T$ Choice Variable (Firm) Length of the free trial period, $T\in\mathbb{N}$
$P$ Choice Variable (Firm) Post-trial automatic renewal price, $P\in(0,1]$
Consumer Choice and Attention
$q$ Choice Variable (Consumer) Probability of successfully remembering to cancel
$\tau(T)$ Outcome Variable Effective attention sensitivity, $\tau(T)=\tau_{0}/(1+\beta T)$
$q^{*}$ Outcome Variable Optimal cancellation probability chosen by the consumer
Key Aggregate Outcomes
$\Pi(T,P)$ Outcome Variable Firm’s total profit
$\text{IR}(T,P)$ Outcome Variable Inattentive Revenue: profit from users who forget to cancel
$U(T,P)$ Outcome Variable Ex-ante expected consumer utility from the contract
$\text{IR-Slack}(T,P)$ Outcome Variable Marginal harm to consumer utility from increasing $T$
Refinement Assumption Parameters
$\varepsilon$ Parameter (A1) Constant price elasticity of "happy subscriber" demand
$\kappa$ Parameter (A1) Scaling constant for iso-elastic demand
$h(v)$ Property (A2) Hazard rate of the valuation distribution, f(v)/(1F(v))f(v)/(1-F(v))
Extension: Paid Trials (Section 6)
$P_{0}$ Choice Variable (Firm) Upfront introductory price for the trial period
$\eta(P_{0})$ Function Fraction of consumers who sign up at intro price $P_{0}$
$\varepsilon_{0}(P_{0})$ Outcome Variable Elasticity of the sign-up rate with respect to $P_{0}$
$P^{\text{aug}}$ Outcome Variable Total expected post-trial profit per subscriber
$\theta$ Parameter (A3) Constant elasticity of sign-up demand
@@
TimeStage 0Stage 1Stage 2Stage 3 Firm commits to $(T,P)$ Consumers observe $v$, decide to sign up Trial period of length $T$ Automatic renewal at price $P$ Case $v\geq P$ Renews: payoff vPv-P, firm revenue $P$ Case $v<P$ Try to cancel: succeeds w.p. $q^{*}(P,\tau(T))$, fails w.p. $1-q^{*}$
@@ 2. MA-2
    Information Cost (Shannon): The cognitive cost to achieve a cancellation probability of $q\in(0,1)$ is given by $C(q;\tau)=\frac{1}{\tau}\mathcal{I}(q)$, where $\mathcal{I}(q)=q\ln q+(1-q)\ln(1-q)$ is the Shannon Mutual Information function and $\tau>0$ is the consumer’s attention sensitivity. This widely used formulation, grounded in information theory and pioneered in economics by Sims (2003), provides a rigorous micro-foundation for costly information processing. It has been successfully applied in numerous contexts, from macroeconomics to marketing, and its key feature is the convexity of the cost function (Matějka and McKay, 2015). This convexity captures the intuitive idea of increasing marginal costs to achieving perfect certainty: it is relatively easy to improve one’s chances of remembering from 50% to 60%, but it requires immense effort to improve from 99% to 100%.
@@ 3. MA-3
    Memory Decay: The consumer’s attention sensitivity $\tau$ is a decreasing function of the trial length $T$, given by the specific functional form $\tau(T)=\frac{\tau_{0}}{1+\beta T}$, with $\tau_{0}>0$ and $\beta>0$. Here, $\tau_{0}$ represents the consumer’s baseline attention for an immediate task, which can be seen as an individual trait (e.g., being naturally organized) or as being influenced by the technological environment (e.g., access to sophisticated calendar apps). The parameter $\beta$ captures the rate of attention decay over time. This assumption is grounded in the psychology framework of prospective memory, which documents that the ability to remember to perform a delayed intention decays over time (McDaniel and Einstein, 2000; Kliegel et al., 2002). The hyperbolic form is chosen for its tractability and its ability to elegantly nest the standard full-attention model (when $\beta=0$) [^2] Hyperbolic decay is also consistent with the common model of delay discounting in behavioral economics (Ainslie, 1991), or the neuroscience finding about animal integration of past experiences Danskin et al. (2023).. When $\beta>0$, the trial length $T$ becomes a strategic instrument for the firm to directly and predictably influence the cognitive friction faced by its customers. In Appendix F, we provide a robustness analysis of alternative functional forms of memory decay.
@@ 3.3 Refinement Assumptions for Tractability
Assumption 1 (A1: Isoelastic Happy-Subscriber Demand).
@@
Streaming services provide a compelling illustration of isoelastic tails. For example, after Netflix’s January 2022 price hike from \$13.99 to \$15.49 (an 11% increase), its subscriber churn rate increased from 2.3% to 3.3%, implying a demand elasticity of approximately 0.4.333Data from Antenna Research, as reported in industry analyses of Netflix’s Q1-2022 performance. Similarly, Spotify’s Premium tiers have exhibited monthly churn in the 1.5% to 2.0% range across multiple price points and markets, consistent with an inelastic demand response. [^4] See, for example, Inderes Equity Research, “Spotify: churn and pricing” (2023). With $\varepsilon<1$ in these prominent cases, our model’s prediction that policies like "click-to-cancel" will raise renewal prices modestly while sharply shortening trial lengths becomes directly relevant.
@@
The valuation distribution $F$ has a positive mass of low-value users ($F(\underline{v})>0$ for some $\underline{v}>0$) and a weakly increasing hazard rate, h(v)=f(v)/(1F(v))h(v)=f(v)/(1-F(v)), on the interval $[0,1]$.
@@
This assumption serves two crucial roles. First, the existence of a low-value segment ensures that there is always a pool of consumers who need to cancel, making the cancellation problem economically relevant. It also ensures that the consumer participation constraint will eventually bind as the trial length grows and the expected costs of managing the subscription become prohibitive for these low-value types. Second, the Increasing Failure Rate (IFR) property is a standard regularity condition in contract theory and industrial organization that guarantees the firm’s profit function is well-behaved (specifically, concave in price), guaranteeing a unique optimal solution (Lariviere and Porteus, 2001). It ensures that the marginal revenue from "happy" subscribers diminishes sufficiently quickly, which is necessary for a well-defined optimum. This assumption is weaker than assuming a specific parametric family; it requires only that the “conditional dropout” rate $h(v)$ does not decline.
@@
3.4 Information–Cost Paradigm (Matějka and McKay, 2015)
To model the consumer’s cancellation decision, we employ the rational inattention (RI) framework based on Shannon entropy, as formalized for discrete choice by Matějka and McKay (2015). This approach posits that processing information to move from a state of uncertainty to a state of certainty is cognitively costly.
Shannon entropy and mutual information.
For a binary random variable that takes value $C\in\{\text{Cancel},\text{No‐Cancel}\}$ with probabilities (q,1q)(q,1-q), the Shannon entropy (in nats) is
$H_{\mathrm{bin}}(q)\;=\;-\Bigl[q\ln q+(1-q)\ln(1-q)\Bigr]\;\;\geq 0,\qquad q\in(0,1).$
Assuming the customer starts from an uninformative prior over the cancellation $p=(\tfrac{1}{2},\tfrac{1}{2})$ [^5] In Appendix F, we show that our results hold in the case of informative priors. The uninformative prior assumption is for brevity., following the Rational Inattention framework (Sims, 2003), we define the "amount" of information that has been used by an customer in making the optimal decision as the mutual information processed by them, or equivalently the Kullback–Leibler divergence of (q,1q)(q,1-q) from $p$, is
$\mathcal{I}(q\parallel p)$ $=q\ln\!\Bigl(\tfrac{q}{1/2}\Bigr)+(1-q)\ln\!\Bigl(\tfrac{1-q}{1/2}\Bigr)$
$=q\ln q+(1-q)\ln(1-q)+\ln 2$ (1)
$=-H_{\mathrm{bin}}(q)\;+\;\ln 2.$
It satisfies $\mathcal{I}(q\parallel p)\geq 0$ with equality at $q=\tfrac{1}{2}$ (no attention).
Definition 1 (Shannon mutual–information cost).
Given an attention sensitivity $\tau>0$, the cognitive cost of implementing reminder probability $q$ is
$C(q;\tau)\;=\;\frac{1}{\tau}\,\mathcal{I}(q\parallel\tfrac{1}{2}).$
Higher $\tau$ makes information processing cheaper. [^6] Here, we use $\tau$ to denote sensitivity, since that is a main focus of our model. Readers familiar with RI framework will realize this is the reciprocal of $\lambda$, the unit cost of attention, in Matějka and McKay (2015).
Because $\ln 2/\tau$ is a constant w.r.t. $q$, we may drop it without affecting any optimization. Accordingly, throughout the proofs we use the algebraically lighter equivalent
$C(q;\tau)=\frac{1}{\tau}\Bigl[q\ln q+(1-q)\ln(1-q)\Bigr].$
The Shannon Mutual information function has several properties that make it a compelling measure of cognitive friction. The parameter $\tau$ is central, acting as an inverse measure of the "price of attention." A consumer with a high $\tau$ may be naturally well-organized, have access to better reminder technology (e.g., sophisticated calendar apps), or face a simpler, less onerous cancellation task; for them, achieving certainty is relatively cheap. The cost function $C(q;\tau)$ itself is convex in $q$, capturing the crucial economic intuition of increasing marginal effort. For instance, moving from complete uncertainty ($q=0.5$) to moderate certainty ($q=0.6$) requires less cognitive effort than moving from high certainty ($q=0.99$) to near-perfect certainty ($q=0.999$). This reflects the real-world difficulty and escalating mental burden of eliminating all possibility of error. Furthermore, the cost is minimized and equal to zero only when no new information is processed and the consumer remains at their uninformative prior ($q=0.5$). This specific functional form is not only tractable, i.e., leading directly to the widely-used logit choice model, but is also deeply rooted in the information-theoretic foundations of the RI literature (Sims, 2003), providing a disciplined micro-foundation for modeling bounded rationality.
@@ 3.5 Consumer Monitoring Problem
To model the cancellation decision of a consumer with $v<P$, we employ the rational inattention framework. The consumer faces a binary choice: ‘Cancel‘ or ‘Do Not Cancel‘. The ‘Cancel‘ action is optimal and yields a payoff of 0. The ‘Do Not Cancel‘ action is a mistake that results in a payment of $P$. The consumer can reduce the probability of making this mistake by devoting cognitive resources (e.g., setting reminders, mentally rehearsing the task) to monitoring the deadline. Let $q\in[0,1]$ be the probability that the consumer successfully remembers to cancel. Achieving a high value of $q$ is cognitively costly. The consumer decision problem is
$\min_{q\in[0,1]}\;\;\Bigl[(1-q)P\Bigr]\;+\;\frac{1}{\tau(T)}\,\mathcal{I}\!\bigl(q\parallel\tfrac{1}{2}\bigr),$
or, dropping the constant $\ln 2/\tau(T)$, this reduces to,
$\min_{q\in[0,1]}\;\;(1-q)P+\frac{1}{\tau(T)}\Bigl[q\ln q+(1-q)\ln(1-q)\Bigr].$
Lemma 1 (Optimal monitoring (RI logit)).
The objective in (3.5) is strictly convex on $(0,1)$ and attains its unique minimum at
$q^{*}(P,\tau)\;=\;\frac{1}{1+\exp(-\tau P)}.$
Moreover, the cancellation probability $q^{*}$ is increasing in price $P$ and attention $\tau$, but decreasing in trial length $T$:
$\frac{\partial q^{*}}{\partial P}=\tau q^{*}(1-q^{*})>0,\quad\frac{\partial q^{*}}{\partial\tau}=P\,q^{*}(1-q^{*})>0,\quad\frac{\partial q^{*}}{\partial T}=-\beta\tau_{0}P\frac{q^{*}(1-q^{*})}{(1+\beta T)^{2}}<0.$
Sketch.
Take the derivative of (3.5): $\partial_{q}L=-P+\tau^{-1}\ln\!\bigl(q/(1-q)\bigr).$ Setting it to zero yields $\ln\!\bigl(q/(1-q)\bigr)=\tau P$, i.e. the stated logit. Strict convexity follows because $\partial_{qq}L=\tau^{-1}/[q(1-q)]>0$. The comparative statics are direct differentiations, and $\partial_{T}q^{*}$ uses the chain rule with $d\tau/dT=-\beta\tau_{0}/(1+\beta T)^{2}$. ∎
The results of this lemma are intuitive. Consumers will try harder to remember to cancel (i.e., choose a higher $q^{*}$) when the financial stakes are higher (a larger $P$) and when paying attention is cognitively cheaper (a higher $\tau$). The result that $q^{*}$ decreases with trial length $T$ is a direct consequence of our memory decay assumption (MA-3) and forms the central behavioral mechanism of our model. The logistic functional form is a hallmark of the Shannon entropy cost structure and provides an analytically tractable foundation for the rest of our analysis.